OpenStudy (anonymous):

Let F(x)=int_{0}^{x}e^(-3t^4) dt Find the MacLaurin polynomial of degree 5 for F(x). Use this polynomial to estimate the value of int_{0}^{0.11}e^(-3x^4) dx

OpenStudy (anonymous):

\[F(x)=\int\limits_{0}^{x}e^(-3t^4) dt\]

OpenStudy (anonymous):

i put 1-3x^4+3x^8/2!-3x^12/3!+3x^16/4! and it is not correct

OpenStudy (anonymous):

\[\int\limits_{0}^{0.11}e^(-3x^4) dx\]

OpenStudy (anonymous):

First let \[ h(x) = e^{-3 x^4} \] The MacLaurin polynmoial of Degree 4 of h at zero is \[1- 3 x^4\] Now the Maclaurin polynomial of degree 5 is \[ g(x)=\int_0^x (1-3 t^4) dt = x - \frac {3 x^5} 5 \] and \[ g[.11] = .10999\]

OpenStudy (anonymous):

i put 0.10999 and it is not correct

OpenStudy (anonymous):

Give 4 or 5 significant digits

OpenStudy (anonymous):

they are correct!!thanks!!